# Friedmann equations in the Codazzi parametrization of Cotton and extended theories of gravity and the Dark Sector

### 05 Dec 2023 v1

@@ V Fixing the dark sector
V.1 Chevallier-Polarski-Linder (CPL) model
@@
V.2 Jassal-Bagla-Padmanabhan (JPB) model
@@
VI Conclusions
Cotton gravity offers a simple setting to reproduce the Friedmann equations of well known extended theories. In all cases the dark sector arising from geometry is described by a Codazzi tensor with the proper choice of a single function. We also showed that the recently proposed conformal Killing gravity is absorbed in Cotton gravity at least for cosmological FRW space-times. The dark sector may be fixed requesting an EoS: this can accomodate in a unified description the best known redshift dependent models.

### 18 Dec 2023 v2

@@ VI Fixing the dark sector
VI.1 Chevallier-Polarski-Linder (CPL) model
@@
VI.2 Jassal-Bagla-Padmanabhan (JPB) model
@@
VII Conclusions
Cotton gravity offers a simple setting to reproduce the Friedmann equations of well known extended theories. In all cases the dark sector arising from geometry is described by a Codazzi tensor with the proper choice of a single function. We also showed that the recently proposed conformal Killing gravity is absorbed in Cotton gravity at least for cosmological FRW space-times; this is also true for Mimetic gravity. The dark sector may be fixed requesting an EoS: this can accomodate in a unified description the best known redshift dependent models.